# High-Capacity Silicon Microdisk Transmitter

### v1

@@ Introduction
The microdisk transmitter chip is designed to support both high parallelism SDM and high per-fiber bandwidth WDM. WDM requires each microring modulating at a distinct dedicated wavelength. On the other hand, fabrication-induced variations in resonance frequency is a major issue in integrated photonic resonant devices and this tolerance cannot be solely achieved through optimized fabrication. While traditional thermal tuning methods are used to compensate for these variations, they suffer from high power consumption. To retune the system to dedicated resonance, we developed a nonvolatile resonance tunning setup based on the laser trimming. Specifically, our trimming method worked for ultra-compact resonant device with top SiO2 cladding and metal wires by changing the top cladding compaction, enabling automated in-situ, close-looped, non-reversal, low-loss, and picometer-precision permanent wavelength tuning, significantly enhancing the controllability compared to previous methods.
@@
To accomplish the laser-based permanent shifting of ring resonances, we introduce an additional free space path to controllably inject a trimming laser onto each ring. This path includes a green laser and critically, a two axis galvanometers for controllable beam focusing and routing (Thorlabs GSV202). This path is merged into the existing chip characterization paths by using an angled half reflective mirror before the objective. We discretize this path into 360 steps around the circle. Every iteration step along the circle takes $\sim 1\leavevmode\nobreak\ \mathrm{s}$, the trimmed spectrum was recorded after each half circle, and summarized in Fig. 4b. Due to the strong mode confinement of silicon waveguide, current trimming mechanism has a small wavelength tuning range, and a minimum wavelength red shift of 100 pm was observed.
Fig. 4c shows the wavelength distribution of the foundry fabricated microdisk modulators, these modulators are designed to have the same resonant wavelength at 1537.4 nm and the actual measured transmission spectrum is separated from 1536.5 nm to 1539 nm. We select the channel with resonances close to the DWDM ITU wavelength, and use the trimming setup to perform automated precisely laser annealing, after trimming, these channels are evenly distributed ($25\leavevmode\nobreak\ \text{GHz}$ spacing) within one FSR.
The OMA, bandwidth and eye-diagram of the laser-trimmed DWDM modulators were experimentally characterized, revealing no discernible degradation after trimming. To mitigate inter-channel crosstalk, channels 1, 3, 5, 7, and 9 were specifically selected for DWDM transmission experiments. The implemented DWDM system features a standard wavelength channel spacing of 50 GHz, with each selected channel independently modulated at a data rate of $20\leavevmode\nobreak\ \mathrm{Gbt/s}$. Consequently, the chip support a fully passive, 5-channel DWDM system with per-fiber capacity of $100\leavevmode\nobreak\ \mathrm{Gbt/s}$. These results underscore the robust performance and scalability of the proposed trimmed EOIM-based transmitter, highlighting its potential as a key component in high-capacity integrated photonic communication networks.
@@
While the current wavelength trimming range is limited, this trimming technique offers two significant practical benefits. The first application is reducing the overall energy consumption in WDM transmission systems. In typical WDM systems, in addition to the dynamic modulation energy, microring resonators require continuous thermal energy supply to maintain the wavelength alignment. By employing the current trimming method, the energy consumption required for tuning and maintaining the resonant wavelength can be effectively reduced by 33 %.
@@
Another important application is reducing the required redundancy of the microring resonators to achieve fully passive WDM transmission. In temperature-sensitive applications, such as cryogenic quantum platforms where active thermal tuning may influence the experimental results, multiple redundant resonators are typically fabricated for each target wavelength to statistically ensure at least one ring naturally aligns to the desired resonance. By employing laser trimming, the necessary redundancy was reduced, effectively halving the number of redundant microrings needed per wavelength to reliably achieve the targeted fully passive WDM operation, as illustrated in Fig. 4f.
@@
Integrated SiPh chips provide a promising platform to meet the escalating demands for high-speed and energy-efficient data movement in modern data centers, particularly for AI applications. Here, we realize this promise by demonstrating a large-scale SiPh transmitter comprising an array of 64 EOIMs. Fabricated on a 300-mm wafer via a fully CMOS-compatible process, each EOIM, based on a compact microdisk resonator with an 8-µm diameter, achieves a high electro-optic tuning efficiency of $90\leavevmode\nobreak\ \mathrm{pm\,V^{-1}}$, a 3-dB electro-optic bandwidth of $20\leavevmode\nobreak\ \mathrm{GHz}$, an ER of $9.8\leavevmode\nobreak\ \mathrm{dB}$, an IL of $2.6\leavevmode\nobreak\ \mathrm{dB}$, and an OMA of 0.49 at a $2\leavevmode\nobreak\ \mathrm{V}$ swing voltage. Moreover, the microdisk modulator exhibits a low $V_{\pi}\cdot L$ of $3.57\leavevmode\nobreak\ \mathrm{V\cdot mm}$. These characteristics enable high-precision, energy-efficient dynamic modulation at $29\leavevmode\nobreak\ \mathrm{fJ/bit}$, supporting data rates up to $20\leavevmode\nobreak\ \mathrm{Gbt/s}$ per channel.
The demonstrated transmitter achieves a data transmission capacity exceeding $1.2\leavevmode\nobreak\ \mathrm{Tbt/s}$, which was limited by our characterization equipment. By utilizing advanced, high-speed AWG sources, modulation speeds of $30\leavevmode\nobreak\ \mathrm{Gbt/s}$ per channel are achievable, enabling scalability to a capacity of $1.8\leavevmode\nobreak\ \mathrm{Tbt/s}$.
Furthermore, we demonstrated that fabrication-induced resonance-frequency variations in the microdisk modulators can be mitigated through an in-situ automated, clos-looped, permanent laser-trimming technique with picometer-scale precision. Utilizing our customized, nonvolatile laser-trimming platform, a passive 5-channel DWDM transmitter system is demonstrated. This trimming method introduces negligible optical losses and preserves the high modulation performance post-trimming. The current trimming technology reduces 33 % thermal energy consumption needed to tune the wavelength. When combined with emerging integrated chip-scale optical frequency comb sources, our scalable DWDM architecture paves the way toward ultra-high-density silicon photonic interconnects capable of supporting massively parallel optical communications. There are no fundamental limits to scaling our approach to thousands of channels in such an architecture. The CMOS compatibility of our fabrication process also enables us to manufacture our photonic architectures directly on top of a high-voltage CMOS driver.
@@ methods
Photonic chip design and fabrication
The active silicon photonic circuits were designed using verified process design kit (PDK) components provided by the foundry. The AIM Integrated PDK used for the transmitter includes the microdisk EOIM, thermal optical phase shifter, optical beam splitter, edge couplers, and vertical optical grating emitter. The mask layout was prepared in GDS Factory. Fabrication was performed as part of the AIM Photonics 300-mm MPW service. The AIM process uses a 220-nm-thick silicon waveguide layer with a 2-μm-thick buried oxide. Additionally, the AIM process and PDK use Si3N4 layers for devices such as inverse-tapered edge couplers and silicon-to-Si3N4 layers escalators. More information regarding the AIM process and PDK is available in [46].
@@
Tuning of the trimming spot is accomplished by scanning the galvanometers angles and makes the first-order beam touted across the waveguide in a dithered circular pattern. We discretize this path into 360 steps around the circle. Every iteration step along the circle takes $\sim 1\leavevmode\nobreak\ \mathrm{s}$. Once the laser intensity and the galvanometers scanning parameters are set, we leave the trimming light constantly on.
After the trimming, wavelength red-shifted is observed, however, due to the strong mode confinement of the silicon waveguide, the laser trimming induced SiO2 compaction change has low influencial for the mode effective index change. So, to target the devices for a specific wavelength we select a chip with resonances close to the desired wavelength. This minimizes the required trimming power/duration. In addition, we observe no significant change in the extracted full width at half maximum given the data in Figure 9. We measure the laser-trimmed devices again after roughly two months of ambient storage and observe no relaxation in trimming. No damage was observed in the microdisk, the E/O modulation and thermal phase shifting still operate well after the laser trimming.
It needed to be point out that the current trimming technology is achieved through changing the SiO2 cladding compaction, the trimming efficiency is related to the SiO2 cladding quality. Using the same trimming methods, we achieve over 0.4 nm wavelength red tunning on some device, and the 0.1 nm wavelength red tuning we used in the manuscript is the minimum wavelength red tunning we obtained among different devices.
@@ Acknowledgements
The authors gratefully acknowledge the slmsuite team for their contributions to the trimming algorithm and code development, and we thank Prof. Karl Berggren (MIT RLE Quantum Nanostructures and Nanofabrication Group) for generously providing access to the vector network analyzer.
@@
Demonstration and Non-volatile Trimming of a High Parallelism,
@@ II.1 Subsection 1: ER, IL, and OMA
The devices exhibit a largest ER of 9.8dB9.8\ $\text${dB} and an IL of 2.6dB2.6\ $\text${dB}. The corresponding average OMA is 0.49.
@@ II.3 Electrical response ($S_{11}$) of the lumped micro-disk and circuit parameter extraction
$S_{11}(\omega)=\dfrac{Z_{\mathrm{in}}(\omega)-Z_{0}}{Z_{\mathrm{in}}(\omega)+Z_{0}},\qquad Z_{0}=50\leavevmode\nobreak\ \Omega.$ (8)
A calibrated VNA measures the complex $S_{11}^{\text{meas}}(\omega)$ from $50\leavevmode\nobreak\ \mathrm{MHz}$ to $40.5\leavevmode\nobreak\ \mathrm{GHz}$ using a GS probe. The substrate contact is left floating so that the entire pad model appears between signal and ground. Although all five parameters can be obtained from a single non-linear fit, we follow the procedure of [16, 17]: the pad resistance is fixed to the software simulation value while the remaining parameters are fitted from the measurements. Figure 7 compares the measured and fitted $S_{11}$; the resulting lumped parameters are summarized in Table 1.
@@ III.1 Subsection 1: Laser annealing
The laser annealing was first verified in a SiN microring E/O modulator, the microring is covered with top thick cladding, through the laser annealing, the cladding bonding angle changes, this will increase the cladding compaction, and increase the refractive index. During the annealing process, we capture the leakage intensity from the microring directly, The measured leakage intensity versus time is shown in Figure 8, as clearly shown, the microring started with off resonance, through the trimming, the microring gradually becomes to in resonance state, and we can see a clear ’ring shape’ leakage, further trimming the microring will make it becomes off resonance again, the leakage intensity gradually increase, when trimmed to the in resonance state, the leakage intensity reach to the highest value, and then gradually decrease with the trimming laser on.
@@
In the silicon micro-disk the out-of-plane scatter port is masked by the integrated heater, so the resonance shift is monitored at the grating-coupler output. A laser spot with a $1\leavevmode\nobreak\ \upmu\text{m}$ radius ($\approx 3\leavevmode\nobreak\ \upmu\text{m}^{2}$) is used for trimming; its power density is $\sim\!1\leavevmode\nobreak\ \text{mW}\,\upmu\text{m}^{-2}$, well above the SiO2 densification threshold. Because the optical mode in silicon is much more tightly confined than in SiN, its evanescent overlap with the cladding—and hence the index change produced by compaction—is smaller. Consequently, the cavity resonance shift is between $\sim\!100\leavevmode\nobreak\ \text{pm}$ to $\sim\!400\leavevmode\nobreak\ \text{pm}$, changes among different microdisk due to the SiO2 quality different of among different devices.
III.2 Subsection 2: Thermal annealing
@@
Thermal annealing is another widely used post-fabrication trimming technique for resonant silicon photonic devices [43, 42]. By holding the chip at several hundred degree, the SiO2 cladding densifies and its refractive index increases, permanently shifting the cavity resonance. Furnace exposures of $\sim 10\text{\textendash}15\leavevmode\nobreak\ \mathrm{min}$ at $400\text{\textendash}500^{\circ}\mathrm{C}$ have been shown to tune passive SiN microrings by more than one free-spectral range with negligible excess loss [43].
@@
Limitation—high temperature. The high process temperature that makes annealing effective for passive structures also limits its applicability to active components. Above $\sim 300^{\circ}\mathrm{C}$ aluminium or Titanium heaters can reflow, dopant profiles in p–n junctions smear out, silicide contacts degrade, metal–oxide capacitors, and p-n diode suffer irreversible damage. Consequently furnace annealing is unsuitable for modulators, photodetectors, or devices that incorporate heterogeneous III–V stacks, where even brief high-temperature exposure can compromise speed or reliability. These constraints motivate low-temperature, highly localized alternatives—such as the laser-induced cladding compaction demonstrated in this work—which trim individual resonators without subjecting the rest of the photonic circuit to damaging thermal budgets. We also implement the thermal annealing method in the current chip, we set the TiN thermal heater driving voltage to 8-V, which makes a 12 nm thermal wavelength red shift, through the silicon thermal modulation efficiency the temperature change within the silicon waveguide is about $250^{\circ}\mathrm{C}$, turning off the heater driving voltage, we find a 200 pm wavelength red shift. To avoid the potential damage to the modulator, we didn’t use this thermal annealing method during the experiment. The issue of the thermal trimmed wavelength shifting is the long time stability, we monitor the resonant wavelength of the thermal trimmed device, and find the resonant wavelength gradually returned to the initial state after 2 weeks, which has been found in other journals [58, 59].
III.3 Subsection 3: Silicon Oxidation

### v2

@@ Introduction
The microdisk transmitter chip is designed to support both high parallelism SDM and high bandwidth density WDM. WDM requires each microdisk modulating at a distinct dedicated wavelength. On the other hand, fabrication-induced variations in resonance frequency is a major issue in integrated photonic resonant devices and this tolerance cannot be solely achieved through optimized fabrication. While traditional thermal tuning methods are used to compensate for these variations, they suffer from high power consumption. To retune the system to dedicated target resonance, we developed a nonvolatile resonance tunning setup based on the laser trimming. Specifically, our trimming method worked for ultra-compact microdisk with top SiO2 cladding and metal embedding by changing the top cladding compaction, enabling automated in-situ, close-looped, non-reversal, low-loss, and picometer-precision permanent wavelength tuning, significantly enhancing the controllability compared to previous methods.
@@
To accomplish the laser-based permanent shifting of microdisk resonances, we introduce an additional free space path to controllably inject a trimming laser onto each ring. This path includes a green laser and critically, a two axis galvanometers for controllable beam focusing and routing (Thorlabs GSV202). This path is merged into the existing chip characterization paths by using an angled half reflective mirror before the objective. We discretize this path into 360 steps around the circle. Every iteration step along the circle takes $\sim 1\penalty 10000\ \mathrm{s}$, the trimmed spectrum was recorded after each half circle, and summarized in Fig. 4b. Due to the strong mode confinement of vertical silicon p-n junction, current trimming mechanism has a small wavelength tuning range from 100 pm to 380 pm.
@@
While the current wavelength trimming range is limited, this trimming technique offers two significant practical applications. The first application is reducing the overall energy consumption in WDM transmission systems to maintain the resonant wavelength. By employing the current trimming method, the energy consumption required for tuning and maintaining the target wavelength can be effectively reduced by 33 %, as illustrated in Fig. 4c.
@@
Another important application is reducing the required design redundancy of the microdisk resonators. In temperature-sensitive applications, such as label-free biosensing (e.g., microring sensors in liquid) [46, 47] and cryogenic quantum platforms [48], the active thermal tuning will influence the experimental results, multiple redundant resonators are typically fabricated for a target wavelength to statistically ensure at least one ring naturally aligns to the desired resonance. We set the design redundancy as the cost function, and the laser trimming was employed to reduce it. By employing laser trimming, the necessary redundancy was reduced, effectively halving cost function, as illustrated in Fig. 4d.
Fully passive DWDM using the Microdisk Transmitter
Fig. 5a shows the wavelength distribution of the foundry fabricated microdisk modulators, these modulators are designed to have the same target resonant wavelength at 1537.2 nm. The actual measured transmission spectrum is separated from 1536.5 nm to 1539 nm. Combined with the non-volatile trimming, this initial distribution can be employed to build the high bandwidth WDM transmission link. We select the channel with resonances close to the DWDM ITU wavelength, and use the trimming setup to perform automated precisely laser annealing, after trimming, these channels are evenly distributed ($25\penalty 10000\ \text{GHz}$ spacing) within one FSR.
The OMA, bandwidth and eye-diagram of the laser-trimmed DWDM modulators were experimentally characterized, revealing no discernible degradation after trimming. To mitigate inter-channel crosstalk, channels 1, 3, 5, 7, and 9 were specifically selected for DWDM transmission experiments. The implemented DWDM system features a standard wavelength channel spacing of 50 GHz, with each selected channel independently modulated at a data rate of $20\penalty 10000\ \mathrm{Gbt/s}$. Consequently, the chip support a fully passive, 5-channel DWDM system with per-fiber capacity of $100\penalty 10000\ \mathrm{Gbt/s}$. These results underscore the robust performance and scalability of the proposed trimmed EOIM-based transmitter, highlighting its potential as a key component in high-capacity integrated photonic communication networks.
@@
Integrated SiPh chips provide a promising platform to meet the escalating demands for high-speed and energy-efficient data movement in modern data centers, particularly for AI applications. Here, we realize this promise by demonstrating a large-scale SiPh transmitter comprising an array of 64 EOIMs. Fabricated on a 300-mm wafer via a fully CMOS-compatible process, each EOIM, based on a compact microdisk resonator with an 8-µm diameter, achieves a high electro-optic tuning efficiency of $90\penalty 10000\ \mathrm{pm\,V^{-1}}$, a 3-dB electro-optic bandwidth of $20\penalty 10000\ \mathrm{GHz}$, an ER of $9.8\penalty 10000\ \mathrm{dB}$, an IL of $2.6\penalty 10000\ \mathrm{dB}$, and an OMA of -3.2 dB relative to input optical power at a $2\penalty 10000\ \mathrm{V}$ swing voltage. Moreover, the microdisk modulator exhibits a low $V_{\pi}\cdot L$ of $3.57\penalty 10000\ \mathrm{V\cdot mm}$. These characteristics enable high-precision, energy-efficient dynamic modulation at $29\penalty 10000\ \mathrm{fJ/bit}$, supporting data rates up to $20\penalty 10000\ \mathrm{Gbt/s}$ per channel.
The demonstrated transmitter achieves a data transmission capacity of $1.28\penalty 10000\ \mathrm{Tbt/s}$, which was limited by our characterization equipment. By utilizing advanced, high-speed AWG sources, modulation speeds of $30\penalty 10000\ \mathrm{Gbt/s}$ per channel are achievable, enabling scalability to a capacity of $1.8\penalty 10000\ \mathrm{Tbt/s}$ [17].
Furthermore, we demonstrated that fabrication-induced resonance-frequency variations in the microdisk modulators can be mitigated through an in-situ automated, clos-looped, permanent laser-trimming technique with picometer-scale precision. Utilizing our customized, nonvolatile laser-trimming platform, a passive 5-channel DWDM transmitter system is demonstrated. This trimming method introduces negligible optical losses and preserves the high modulation performance. The current trimming technology reduces 33 % thermal energy consumption needed to tune the foundry microdisk to target resonant wavelength, and effectively halving the number of redundant microdisks needed for temperature sensitive applications where thermal tunning is not allowed. When combined with emerging integrated chip-scale optical frequency comb sources, our scalable DWDM architecture paves the way toward ultra-high-density silicon photonic interconnects capable of supporting massively parallel optical communications.
@@ methods
Tuning of the trimming spot is accomplished by scanning the galvanometers angles and makes the first-order beam touted across the waveguide in a dithered circular pattern. We discretize this path into 360 steps around the circle. Every iteration step along the circle takes $\sim 1\penalty 10000\ \mathrm{s}$. Once the laser intensity and the galvanometers scanning parameters are set, we leave the trimming light constantly on.
After the trimming, wavelength red-shifted is observed, however, due to the strong mode confinement of the silicon waveguide, the laser trimming induced SiO2 compaction change has low influential for the mode effective index change. So, to target the devices for a specific wavelength we select a chip with resonances close to the desired wavelength. This minimizes the required trimming power/duration. In addition, we observe no significant change in the extracted full width at half maximum given the data in Figure 10. We measure the laser-trimmed devices again after roughly two months of ambient storage and observe no relaxation in trimming. No damage was observed in the microdisk, the E/O modulation and thermal phase shifting still operate well after the laser trimming.
It needed to be point out that the current trimming technology is achieved through changing the SiO2 cladding compaction, the trimming efficiency is related to the SiO2 cladding quality. Using the same trimming methods, we achieve over 0.35 nm wavelength red tunning on some device, and the 0.1 nm wavelength red tuning we used in the manuscript is the minimum wavelength red tunning we obtained among different devices.
@@ Acknowledgements
The authors gratefully acknowledge the slmsuite team for their contributions to the trimming algorithm and code development.
@@
Demonstration and Non-volatile Trimming of a Highly Parallel,
@@ II.1 Subsection 1: ER, IL, and OMA
The devices exhibit a largest ER of 9.8dB9.8\ $\text${dB} and an IL of 2.6dB2.6\ $\text${dB}. The corresponding average OMA is -3.2 dB response to the input power in a 2-V voltage swing.
@@ II.3 Electrical response ($S_{11}$) of the lumped micro-disk and circuit parameter extraction
$S_{11}(\omega)=\dfrac{Z_{\mathrm{in}}(\omega)-Z_{0}}{Z_{\mathrm{in}}(\omega)+Z_{0}},\qquad Z_{0}=50\penalty 10000\ \Omega.$ (8)
A calibrated VNA measures the complex $S_{11}^{\text{meas}}(\omega)$ from $50\penalty 10000\ \mathrm{MHz}$ to $40.5\penalty 10000\ \mathrm{GHz}$ using a GS probe. The substrate contact is left floating so that the entire pad model appears between signal and ground. Although all five parameters can be obtained from a single non-linear fit, we follow the procedure of [16, 17]: the pad resistance is fixed to the software simulation value while the remaining parameters are fitted from the measurements. Figure 8 compares the measured and fitted $S_{11}$; the resulting lumped parameters are summarized in Table 1.
@@ III.1 Subsection 1: Laser annealing
The laser annealing was first verified in a SiN microring E/O modulator, the microring is covered with top thick cladding, through the laser annealing, the cladding bonding angle changes, this will increase the cladding compaction, and increase the refractive index. During the annealing process, we capture the leakage intensity from above the microring directly, The measured leakage intensity versus time is shown in Figure 9, as clearly shown, the microring started with off resonance, through the trimming, the microring gradually becomes to in resonance state, and we can see a clear ’ring shape’ leakage, further trimming the microring will make it becomes off resonance again, the leakage intensity gradually increase, when trimmed to the in resonance state, the leakage intensity reach to the highest value, and then gradually decrease with the trimming laser on.
@@
In the silicon micro-disk the out-of-plane scatter port is masked by the integrated heater, so the resonance shift is monitored at the grating-coupler output. A laser spot with a $1\penalty 10000\ \upmu\text{m}$ radius ($\approx 3\penalty 10000\ \upmu\text{m}^{2}$) is used for trimming; its power density is $\sim\!1\penalty 10000\ \text{mW}\,\upmu\text{m}^{-2}$, well above the SiO2 densification threshold. Because the optical mode in silicon is much more tightly confined than in SiN, its evanescent overlap with the cladding—and hence the index change produced by compaction—is smaller. Consequently, the cavity resonance shift is between $\sim\!100\penalty 10000\ \text{pm}$ to $\sim\!380\penalty 10000\ \text{pm}$, changes among different microdisk due to the SiO2 quality different of among different devices.
III.2 Subsection 2: Reduce the thermal tunning energy consumption
The thermal energy required to tune a resonator scales with the wavelength shift needed to reach the target, $\Delta\lambda=\lambda_{\mathrm{tar}}-\lambda$. Because our trimming method provides only a red shift, devices whose initial resonance exceeds the target $(\lambda_{i}>\lambda_{\mathrm{tar}})$ are excluded from analysis.
For microdisks with $\lambda_{i}<\lambda_{\mathrm{tar}}$, we apply a trimming cap $C$ and define the post-trim resonance for device $i$ as
$\lambda_{i}^{\prime}(C)=\min\!\bigl(\lambda_{i}+C,\ \lambda_{\mathrm{tar}}\bigr).$ (9)
We then compute the ensemble-average resonance over all included devices,
$\bar{\lambda}(C)=\langle\lambda_{i}^{\prime}(C)\rangle,$ (10)
and the residual detuning relative to the target,
$\varepsilon(C)=\lambda_{\mathrm{tar}}-\bar{\lambda}(C).$ (11)
Since the thermal energy required after trimming is proportional to the remaining wavelength error, the relative energy saving is,
$S(C)=1-\frac{\varepsilon(C)}{\varepsilon(0)},$ (12)
where $\varepsilon(0)$ is the untrimmed residual detuning computed from the pre-trim ensemble average.
III.3 Subsection 3: Reduce the design redundancy
We model each of the $K$ WDM channels as provisioned with redundancy $r$ (i.e., up to $r$ rings per channel). Assuming each ring independently lands within the target passband with probability $p$, the channel success probability is
$p_{\mathrm{ch}}(p,r)=1-(1-p)^{r}.$ (13)
Hence, the expected number of usable WDM channels is
$\mathbb{E}[N]=K\,p_{\mathrm{ch}}(p,r).$ (14)
For a target of at least $N_{\mathrm{tgt}}$ usable channels in expectation, the minimal redundancy that satisfies $\mathbb{E}[N]\geq N_{\mathrm{tgt}}$ is
$r^{\star}(p;N_{\mathrm{tgt}},K)\;=\;\left\lceil\frac{\ln\!\bigl(1-N_{\mathrm{tgt}}/K\bigr)}{\ln(1-p)}\right\rceil.$ (15)
Trimming increases the per-ring success probability from $p_{\mathrm{pre}}=0.0468$ to $p_{\mathrm{post}}=0.14$. Using (15) with different $K$ number, the required redundancy was calculated, as indicated in Figure 9.
III.4 Subsection 4: Thermal annealing
@@
Thermal annealing is another widely used post-fabrication trimming technique for resonant silicon photonic devices [43, 42]. By holding the chip at several hundred degree, the SiO2 cladding densifies and its refractive index increases, permanently shifting the cavity resonance. Furnace exposures of $\sim 10\text{\textendash}15\penalty 10000\ \mathrm{min}$ at $400\text{\textendash}500^{\circ}\mathrm{C}$ have been shown to tune passive SiN microrings by more than one free-spectral range with negligible excess loss [43].
@@
Limitation—high temperature. The high process temperature that makes annealing effective for passive structures also limits its applicability to active components. Above $\sim 300^{\circ}\mathrm{C}$ aluminium or Titanium heaters can reflow, dopant profiles in p–n junctions smear out, silicide contacts degrade, metal–oxide capacitors, and pnp-n diode suffer irreversible damage. Consequently furnace annealing is unsuitable for modulators, photodetectors, or devices that incorporate heterogeneous III–V stacks, where even brief high-temperature exposure can compromise speed or reliability. These constraints motivate low-temperature, highly localized alternatives—such as the laser-induced cladding compaction demonstrated in this work—which trim individual resonators without subjecting the rest of the photonic circuit to damaging thermal budgets. We also implement the thermal annealing method in the current chip, we set the thermal heater driving voltage to 8-V, the calculated temperature in the silicon waveguide is about $300^{\circ}\mathrm{C}$, turning off the heater driving voltage, we find a 200 pm wavelength red shift. To avoid the potential damage to the modulator, we didn’t use this thermal annealing method during the experiment. The issue of the thermal trimmed wavelength shifting is the long time stability, we monitor the resonant wavelength of the thermal trimmed device, and find the resonant wavelength gradually returned to the initial state after 2 weeks, which has been found in other journals [60, 61].
III.5 Subsection 5: Silicon Oxidation